### v1

@@ 4 Conclusions
*Author contributions A. Mazidi: conceptualization, methodology, software, validation, formal analysis, investigation, data curation, writing – original draft, writing – review and editing, visualization. M. Roshanfar: formal analysis, writing – review and editing. A. Sayadi: methodology, validation, formal analysis, writing – review and editing. J. Dargahi: conceptualization, resources, supervision, writing – review and editing, funding acquisition. J. Barralet: clinical guidance, methodology, writing – review and editing. L. Feldman: clinical guidance, supervision, writing – review and editing. A. Hooshiar: conceptualization, methodology, resources, supervision, writing – review and editing, project administration, funding acquisition.
[…]
@@
*Acknowledgments This research was supported by the Natural Science and Engineering Research Council (NSERC) of Canada through the NSERC CREATE Grant for Innovation at the Cutting Edge (ICE), the Fonds de Recherche du Québec pour la Nature et les Technologies (FRQNT), Concordia University, Research Institute of McGill University Health Centre (RI-MUHC), Montreal General Hospital Foundation (MGHF), and McGill University.
[…]
*Ethics Statement The authors have nothing to report.
[…]
*Conflict of interest The authors declare no conflicts of interest.
[…]
@@
*Data Availability Statement The data that support the findings of this study are available from the corresponding author upon reasonable request.
[…]
@@
*Permission to Reproduce Material From Other Sources All figures, images, and content in this paper are original works created by the authors. No material has been reproduced from other sources.

### v2

@@ 5 Conclusions
A. Mazidi: conceptualization, methodology, software, validation, formal analysis, investigation, data curation, writing, original draft, writing, review and editing, visualization. M. Roshanfar: formal analysis, writing, review and editing. A. Sayadi: methodology, validation, formal analysis, writing, review and editing. J. Dargahi: conceptualization, resources, supervision, writing, review and editing, funding acquisition. J. Barralet: clinical guidance, methodology, writing, review and editing. L. Feldman: clinical guidance, supervision, writing, review and editing. A. Hooshiar: conceptualization, methodology, resources, supervision, writing, review and editing, project administration, funding acquisition.
Acknowledgments
@@
This research was supported by the Natural Science and Engineering Research Council (NSERC) of Canada through the NSERC CREATE Grant for Innovation at the Cutting Edge (ICE), the Fonds de Recherche du Québec pour la Nature et les Technologies (FRQNT), Concordia University, Research Institute of McGill University Health Centre (RI-MUHC), Montreal General Hospital Foundation (MGHF), and McGill University.
Ethics Statement
This study did not involve human participants, human tissue, or animals. All experiments were conducted using phantom models (mannequins with tissue-representative surrogates). Therefore, ethics approval was not required.
Conflict of interest
The authors declare no conflicts of interest.
Data Availability Statement
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Permission to Reproduce Material From Other Sources
@@
All figures, images, and content in this paper are original works created by the authors. No material has been reproduced from other sources.
Author Biography
Aiden Mazidi is a Master of Science candidate at Concordia University, where he serves as a Research Assistant at both McGill University’s Surgical Performance Enhancement and Robotics Center (SuPER) and the Surgical Robotic Laboratory (SRL) at Concordia University. Aiden’s research is focused on force rendering and haptic feedback, with a particular emphasis on RAMIS. His work involves advancing precision surgical interventions through innovative technologies in collaboration with the Surgical Innovation Program at McGill University.
Majid Roshanfar is a Postdoctoral Research Fellow at the Wilfred and Joyce Posluns Centre for Image-Guided Innovation and Therapeutic Intervention (PCIGITI) at The Hospital for Sick Children (SickKids) and the University of Toronto, Canada. He received his Ph.D. in Mechanical Engineering from Concordia University, where his research focused on hybrid-actuated soft robots with stiffness adaptation for surgical applications. His current work focuses on continuum manipulators, magnetic actuation, and force sensing for MIS.
Amir Sayadi is a Ph.D. candidate and researcher at the Surgical Performance Enhancement and Robotics Centre (SuPER) at McGill University, Montreal, Canada. His research focuses on medical robotics, haptic interfaces, and intelligent control systems for minimally invasive and image-guided surgical applications. His work aims to advance the development of robotic platforms that improve surgical precision.
Javad Dargahi received a Ph.D. degree in robotic tactile sensing from Glasgow Caledonian University. He joined the Department of Mechanical and Industrial Engineering, at Concordia University, Montreal, QC, where he is currently a Professor with the Department of Mechanical Engineering. He has co-authored three books in the scope of mechatronics and tactile sensing and has published more than 200 peer-reviewed articles. His research focuses on haptic sensors and feedback for MIS and robotics, as well as micro-manufacturing of sensors and actuators.
Jake Barralet is currently the Vice Chair of Research and Alan Thompson Chair of Surgical Research in the Department of Surgery at McGill University, Montreal, Canada. He also serves as the Scientific Director of the McGill Clinical Innovation Platform and the Associate Director of the Surgical and Interventional Sciences Program at the Research Institute of the McGill University Health Centre. His research focuses on surgical innovation and material-mediated tissue regeneration, with an emphasis on the role of inorganic ions in bone and skin healing.
Liane S. Feldman is the Edward W. Archibald Professor and Chair of Surgery at McGill University and the Surgeon-in-Chief at the McGill University Health Centre. Her clinical focus is on minimally invasive gastrointestinal surgery, and her research aims to improve recovery and outcomes following abdominal operations. She leads a multidisciplinary team implementing evidence-based Enhanced Recovery care plans recognized by Accreditation Canada. She has served as the President of the Society of American Gastrointestinal and Endoscopic Surgeons and has held leadership positions in the Canadian Association of General Surgeons and the American College of Surgeons.
Amir Hooshiar is an Assistant Professor, Edwards Distinguished Scientist, and the Director of Surgical Performance Enhancement and Robotics Centre (SuPER), at the Department of Surgery, McGill University. He specializes in the development of surgical robots and associated technologies such as surgical planning and navigation, sensing and actuation with smart materials, learning-based modelling and control of surgical robots.